OpenStudy (anonymous):

Solve 7 square root of 13 x end root minus 15 square root of 2 x end root plus 2 square root of 13 x end root minus 4 square root of 2 x

OpenStudy (anonymous):

what's an end root?

OpenStudy (anonymous):

Just ignore the end part

OpenStudy (anonymous):

\[7\sqrt{13x}-15\sqrt{2x}+2\sqrt{13x}-4\sqrt{2x},\]\[9\sqrt{13x}-19\sqrt{2x}...\]and this is equal to?

OpenStudy (anonymous):

i guess just simplify?

OpenStudy (anonymous):

but you said solve, so is there an equation?

OpenStudy (anonymous):

Yeah but I don't know how to do it. Do you just add/subtract like you would 2+2 or 2-2?

OpenStudy (anonymous):

just add/subtrac like terms...

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

You can only do that to like-terms. For example, in your case,\[7\sqrt{13x}-15\sqrt{2x}+2\sqrt{13x}-4\sqrt{2x}\]is the same as\[\sqrt{13x}(7+2)+\sqrt{2x}(-15-4).\]Do you see this?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so that answer given by @pre-algebra in the last line is correct

OpenStudy (anonymous):

By the way, your problem cannot be simplified any further than\[\sqrt{x}(9\sqrt{13}-19\sqrt{2})\]

OpenStudy (anonymous):

okay thank you so much